# Assortment Optimization Under History-Dependent Effects

### v2

@@ 2.2 Hardness Result
Our model (HAP) is NP-hard even when the planning horizon is two and the memory length is one.
2.3 Revenue-Ordered Policies and Optimality Structures
A prevalent strategy for solving assortment optimization is to construct RO assortments, which are a family of nested assortments selected according to their revenue-decreasing order. Specifically, RO structures and their variants have been shown to be optimal for solving assortment optimization problems under MNL choice models (Talluri and Van Ryzin 2004), sequential MNL choice models (Gao et al. 2021), and MNL choice models with network effects (Wang and Wang 2017). Inspired by the simplicity and optimality of RO structures, we would like to understand how it works in our problem.
We find that when there are no cross-product and cross-period constraints, if the history-dependent effect is non-negative, that is, $\boldsymbol{\beta}\geq 0$, the optimal assortment in each period is a RO assortment and the assortment planning can be obtained by a sequential revenue-ordered policy. The sequential RO policy is obtained by sequentially selecting the best RO assortment based on the currently updated product utility and maximizing revenue in the current period. This policy is adaptive while ignoring the impact of the current assortment decision on future revenue. Details of the sequential RO policy is shown in Algorithm 1.
Algorithm 1 Sequential revenue-ordered policy of (HAP)
1:  Input: $\mathbf{r},(\beta_{i}^{0})_{i\in[N]},\boldsymbol{\beta},M,N,T$
2:  for all $t=1,2,\dots,T$ do
5:  end for
6:  Output: $\boldsymbol{x}^{t},t\in[T]$
Because there are no cross-product and cross-period constraints, the computation time required to find the optimal RO assortment is $\mathcal{O}(N)$ in each period (Talluri and Van Ryzin 2004, Liu and Van Ryzin 2008). Thus, the total computation time of the sequential RO policy is $\mathcal{O}(NT)$.
Theorem 2.2
Assume that there are no cross-product or cross-period constraints. Then, the sequential-revenue-ordered policy solves model (HAP) if the history-dependent effects are non-negative, that is, $\boldsymbol{\beta}\geq\boldsymbol{0}$.
For the case of negative history-dependent effects, which is the computationally intractable regime of our model, as we argued in Proposition 2.1, the sequential policy fails to solve our model. However, the RO structure still exists and appears in a different form. For an optimal assortment planning of model (HAP), we argue that the set of products that are offered across all periods has a RO structure.
Proposition 2.3
Assume that there are no cross-product or cross-period constraints. Then, there exists an optimal assortment planning $\boldsymbol{x}=(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{T})$ of model (HAP) such that $\bigl{\{}i\in[N]\bigm{|}\sum_{t\in T}x^{t}_{i}\geq 1\bigr{\}}$ is revenue-ordered.
We conclude this section with an example to numerically illustrate that the optimality gap of RO-type policies can be significant for a general instance of (HAP). This motivates us to study the global optimality of (HAP) in Section 3.
Example 2.4
In this example, we compare the revenue of two heuristic policies and the best feasible solution obtained within 3600 seconds for randomly generated instances. The revenue and base utility of an product are randomly generated from uniform distributions with ranges of $[1,10]$ and $[-1,1]$, respectively. The history-dependent effect $\beta_{i}^{m}$ is uniformly sampled from $[-2,-1]$. One of the heuristic policies is the sequential RO-policy. The other is a repeated policy, a non-adaptive policy that chooses an optimal assortment for the first period and repeatedly offers the same assortment for the remaining $T-1$ periods. The relative revenue gap between a myopic policy and the feasible solution is
$G_{\text{rev}}=100\%\times\frac{R_{\mathsf{fea}}-R_{\mathsf{myopic}}}{R_{\mathsf{fea}}},$
where $R_{\mathsf{fea}}$ is the revenue generated by the best feasible solution of (HAP) achieved within 3600s and $R_{\mathsf{myopic}}$ is the revenue generated by one of two myopic policies.
Table 1: Revenue gap of myopic policies
[…]
N T $M=1$ $M=2$ $M=1$ $M=2$
30 5 8.26 14.86 13.84 35.66
30 10 10.21 26.12 15.78 46.41
50 5 7.17 12.48 12.93 32.71
50 10 27.18 22.12 32.34 42.35
70 5 5.14 7.85 10.71 23.05
70 10 6.37 14.45 12.03 29.71
[…]
3 Mixed-Integer Programming Formulation of (HAP)
@@
In this section, we aim to solve the history-dependent assortment problem (HAP) to its global optimality. Due to the combinatorial and nonconvex nature of (HAP), state-of-the-art global optimization solvers such as BARON (Tawarmalani and Sahinidis 2005) and SCIP (Bestuzheva et al. 2021) fail to solve practical-sized instances—see Section 6.1 for computational evidence. Nevertheless, we focus on developing techniques for obtaining mixed-integer programming formulations of (HAP), which will allow us to leverage modern mixed-integer program solvers such as GUROBI and MOSEK. Then, in Section 4, we will use these techniques to find asymptotically optimal cyclic policies.
@@
$\max_{\boldsymbol{x},\boldsymbol{y},\rho}\quad$ $\sum_{i\in[N]}\frac{1}{T}\sum_{t\in[T]}r_{i}y_{i}^{t}$
[…]
@@
$\alpha_{i}(z^{1},\ldots,z^{M}):=\exp\biggl{(}\beta_{i}^{0}+\sum_{k\in[M]}\beta^{k}_{i}\cdot z^{k}\biggr{)}\qquad\operatorname{for}\;(z^{1},\ldots,z^{M})\in\{0,1\}^{M}.$
For a given assortment $\boldsymbol{x}:=(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{T})$, we interpret $\alpha_{i}(x^{t-1},\ldots,x_{i}^{t-M})$ as the attractiveness of product $i$ in period $t$. Therefore, we will henceforth refer to $\alpha_{i}(\cdot)$ as the attraction value function. To handle the discrete nature of the attraction value function, we will use the notion of a continuous extension. Given a function $f:\{0,1\}^{n}\to\mathbbm{R}$, a function $g:[0,1]^{n}\to\mathbbm{R}$ is called a continuous extension of $f(\cdot)$ if $g(\cdot)$ is continuous and $g(\boldsymbol{z})=f(\boldsymbol{z})$ for every $\boldsymbol{z}\in\{0,1\}^{n}$. In addition, if an extension function $g(\cdot)$ is convex (resp. concave), we say that it is a convex (resp. concave) extension. It is common for a discrete function to admit many convex (resp. concave) extensions. A natural idea is to use the tightest convex/concave extension. Since the discrete domain $\{0,1\}^{n}$ is the vertex set of the continuous domain $[0,1]^{n}$, the pointwise largest convex (resp. smallest concave) extension of $f(\cdot)$ is the convex (resp. concave) envelope, denoted as $\operatorname{conv}(f)(\cdot)$ (resp. $\operatorname{conc}(f)(\cdot)$), of $f(\cdot)$ over the box $[0,1]^{n}$ (see Tawarmalani and Sahinidis 2002, Theorem 6). Recall that the convex (resp. concave) envelope of $f(\cdot)$ is referred to the tightest under (resp. over)-estimator of $f$ over $[0,1]^{n}$ (Rockafellar 1997).
3.1 A Lifted Representation of (Choice)
In this subsection, we focus on modelling constraint (Choice). The idea is to lift the original nonconvex constraint (Choice) into a higher-dimensional space where the lifted constraint is dramatically simpler to represent than the original constraint. Specifically, this will allow us to invoke prevalent techniques such as the McCormick envelope (McCormick 1976), the perspective formulation (Günlük and Linderoth 2010), and the Lovász extension (Lovász 1983) to model the lifted nonconvex structures. We elaborate on this in the following Sections 3.1.1 and 3.1.2.
3.1.1 Lifting for Simplicity
@@
$\lambda+\lambda f(\boldsymbol{x})\leq 1,$
where $\lambda\geq 0,\boldsymbol{x}\in\mathbbm{R}^{n}$, and $f(\cdot)$ is a convex function. This constraint can be relaxed into a convex constraint as follows. First, we multiply and divide $\boldsymbol{x}$ by $\lambda$ to obtain
@@
where the first constraint can be linearized using the McCormick envelope McCormick (1976) and the second constraint is convex, since the function $\lambda f(\boldsymbol{u}/\lambda)$ is the perspective function of the convex function $f(\cdot)$ (Rockafellar 1997). Formally, given a function $f:D\to\mathbbm{R}$, where $D$ is a bounded subset of $\mathbbm{R}^{n}$, the perspective function of $f$, denoted as $\operatorname{pers}(f)$, is defined as
@@
Constraint (2a) lifts the nonlinear scaling factor $\rho^{t}x_{i}^{t}$ to $\gamma_{i}^{t}$; constraint (2b) lifts the scaled historical assortment decision $\gamma_{i}^{t}x_{ik}^{t}$ to $z_{ik}^{t}$; and constraint (2c) perspectifies the attraction value function in the space of the newly introduced variables $\gamma_{i}^{t}$ and $\boldsymbol{z}_{i}^{t}$.
@@
$z_{im}^{t}$ $\leq\min\bigl{\{}\gamma_{i}^{t},\rho^{t}_{U}x^{(t-m)}_{i}\bigr{\}}$ $\quad\operatorname{for}\,m\in[M]$ (4)
$z_{im}^{t}$ $\geq\max\bigl{\{}0,\ \rho^{t}_{U}x^{(t-m)}_{i}+\gamma^{t}_{i}-\rho^{t}_{U}\bigr{\}}$ $\quad\operatorname{for}\,m\in[M].$
@@
This is a convex inequality since $\tilde{\alpha}_{i}(\cdot)$ is a convex function and so is its perspective function (Rockafellar 1997). More precisely, this constraint can be represented as an exponential cone constraint (Ben-Tal and Nemirovski 2001) (see more details on obtaining an exponential cone representation in Appendix 9.3):
@@
In the following Section 3.1.2, we show that the concave envelope of $\alpha_{i}(\cdot)$ can be described using the Lovász extension of $\alpha_{i}(\cdot)$ (Lovász 1983).
3.1.2 The Lovász Extension of $\alpha_{i}(\cdot)$
We now derive a closed-form expression of the concave envelope of the attraction value function, which appears in constraint (6). To do this, we exploit the fact that the attraction value function $\alpha_{i}(\cdot)$ is supermodular. Recall that a function $f:\{0,1\}^{n}\to\mathbbm{R}$ is supermodular if
$f(x^{\prime}\vee x^{\prime\prime})+f(x^{\prime}\wedge x^{\prime\prime})\geq f(x^{\prime})+f(x^{\prime\prime})\quad\operatorname{for}x^{\prime},x^{\prime\prime}\in\{0,1\}^{n},$
where $x^{\prime}\vee x^{\prime\prime}$ (resp. $x^{\prime}\wedge x^{\prime\prime}$) is the coordinate-wise maximum (resp. minimum) of $x^{\prime}$ and $x^{\prime\prime}$.
Lemma 3.1
If $\boldsymbol{\beta}\leq 0$ then the attraction value function $\alpha_{i}(\cdot)$ is supermodular over $\{0,1\}^{M}$.
Function $\alpha_{i}(\cdot)$ maps the offering decisions for product $i$ over the past $M$ periods to an attraction value. Given that the history-dependent effects are nonpositive, the more number of times that product $i$ has been offered in the past, the lower its current attraction value. Consequently, increasing the number of past offerings results in smaller changes to the attraction value due to the convexity of the attraction value function. This means that, as product $i$ is offered more in the past, the decay in its attraction value diminishes with each additional offering, indicating a complementary effect (Topkis 1998).
@@ 3.1.2 The Lovász Extension of $\alpha_{i}(\cdot)$
The supermodularity of $\alpha_{i}(\cdot)$ allows us to use the Lovász extension to describe its concave envelope (Lovász 1983, Tawarmalani et al. 2013, Atamtürk and Narayanan 2022). The Lovász extension is an extension of a function $f$ defined on $\{0,1\}^{n}$ to a function $\hat{f}$ defined on $[0,1]^{n}$. For each subset $S$ of $[n]$, let $\chi^{S}\in\{0,1\}^{M}$ be its indicator vector—that is, the $i^{\text{th}}$ coordinate of $\chi^{S}$ is $1$ if and only if $i\in S$. Observe that every vector $\boldsymbol{x}\in[0,1]^{n}$ can be expressed uniquely as $\boldsymbol{x}=\lambda_{1}\chi^{T_{1}}+\lambda_{2}\chi^{T_{2}}+\cdots\lambda_{k}\chi^{T_{k}}$, where $\lambda_{i}\geq 0$ for $i=1,\ldots,k$ and $\sum_{k=0}^{n}\lambda_{k}=1$, and $\emptyset\subseteq T_{0}\subset T_{1}\subset\cdots\subset T_{k}\subseteq[n]$. Thus,
$\hat{f}(\boldsymbol{x})=\lambda_{1}f(\chi^{T_{1}})+\lambda_{2}f(\chi^{T_{2}})+\cdots+\lambda_{k}f(\chi^{T_{k}})$
@@
is a well-defined extension of the function $f$ (called the Lovász extension of $f$) on the continuous domain $[0,1]^{n}$. The Lovász extension serves as a bridge between supermodular functions and concave functions: $f$ is supermodular if and only if its Lovász extension $\hat{f}$ is concave (Lovász 1983). Moreover, when the Lovász extension $\hat{f}$ is concave, it is expressible as the pointwise minimum of affine functions (Tawarmalani et al. 2013). In the following lemma, we show that when these results are scaled with the variable $\gamma^{t}_{i}$, we can obtain an explicit polyhedral description of constraint (6).
Lemma 3.2
Let $\Omega$ be the set of permutations of memory length $\{1,2,\ldots,M\}$. For a permutation $\sigma\in\Omega$, define a collection of nested historical offerings in $\{0,1\}^{M}$ as follows: $\boldsymbol{h}^{\sigma}_{0}=(0,0,\ldots,0)$ and $\boldsymbol{h}^{\sigma}_{k}=\boldsymbol{h}^{\sigma}_{k-1}+\boldsymbol{e}_{\sigma(k)}$ for $k\in[M]$, where $\boldsymbol{e}_{k}$ is the $k^{\text{th}}$ principal vector in $\mathbbm{R}^{M}$. Then, under the presence of constraint (4), constraint (6) is equivalent to the following system of linear inequalities:
@@
$y^{t}_{i}\leq\alpha_{i}(\boldsymbol{h}^{\sigma}_{0})\cdot\gamma^{t}_{i}+\sum_{k\in[M]}\bigl{(}\alpha_{i}(\boldsymbol{h}^{\sigma}_{k})-\alpha_{i}(\boldsymbol{h}^{\sigma}_{k-1})\bigr{)}\cdot z^{t}_{ik}\qquad\operatorname{for}\sigma\in\Omega.$ (7)
3.2 The Final Formulation

### v3

@@ 2.1 Incoporating History-Dependent Effects into Choice Models
(HAP) is NP-hard even when the planning horizon is two, the memory length is one, and history-dependent effects are negative.
2.2 Model Discussion
We outline here the main features of our model, justify our assumptions, and identify relevant scenarios.
It is apparent that our formulation requires a firm to present the same assortment to all customers in each period before the planning horizon begins. In other words, the firm is unable to take advantage of personalized assortments and adaptive policies, which are common practices in online assortment optimization such as Amazon’s recommendation system. However, those practices are not suitable for the case of corporate cafeterias, where it is common to treat employees with the same menu, and it is impractical to dynamically adjust assortments due to the lead time in both procurement and cooking processes (Van Wezel et al. 2006). Observations from our partnered catering company indicate that the corporate cafeteria plans the lunch menus on a weekly basis. Moreover, our model assumption is also reasonable for flash sale channels (i.e., predetermine assortments for several days due to the negotiation time with suppliers).
Our model describes the history-dependent effect through a linear utility function with historical offerings as features. The linear utility function assumption is common in the logit-based choice model to capture impacts such as product attributes (Guadagni and Little 1983) and historical prices (Wang 2018). Moreover, the linear utility function covers multiplication in the attraction value in literature. For example, our model with negative history-dependent effects represents the model in Caro et al. (2014). Our model offers additional flexibility by allowing for positive, negative, and mixed history-dependent effects and offering products in any period.
In the linear utility function, we incorporate the historical offerings to capture the history-dependent effect. The utility function captures the direct effect of historical assortments that are critical decisions at a firm level. Moreover, we only consider the direct history-dependent effect, i.e., the historical offering of a product will only affect the product’s future utility itself. We do not include the cross-product history-dependent effect, i.e., the historical offering of a product might affect the utility of other products, because this effect is usually at a much smaller scale (e.g., Vilcassim et al. 1999, Dubé and Manchanda 2005, Mehta 2007). In our empirical estimation, the cross-product history-dependent effect is approximately one-tenth of the direct effect and is significant only when $M=1$. We believe that our model efficiently captures the primary effect and maintains model tractability.
We manage the history-dependent effects across periods through the memory length $M$, inspired by the patience level from dynamic pricing studies (Popescu and Wu 2007, Liu and Cooper 2015). This patience level indicates how many periods a customer will wait for a price reduction. Consequently, it is logical to assume that customers might retain a limited memory length for a product.
Our model does not include inventory decisions. In cafeteria settings, for instance, procurement choices can be made subsequent to assortment planning. Additionally, inventory rollover is not considered in the study because cooked dishes cannot be kept for the following day due to hygiene standards. To simplify the analysis, this study will concentrate on the assortment planning problem with cross-product and cross-period effects, which is already a challenging problem.
3 Solution Approaches
In this section, we aim to solve the history-dependent assortment problem (HAP) to its global optimality. In Section 3.1, we propose a mixed-integer exponential cone formulation for the general case, and, in Section 3.2, we characterize conditions under which a sequential revenue-ordered policy is optimal.
3.1 Integer Programming Approach
@@
Due to the combinatorial and nonconvex nature of (HAP), state-of-the-art global optimization solvers such as BARON (Tawarmalani and Sahinidis 2005) and SCIP (Bestuzheva et al. 2021) fail to solve practical-sized instances—see Section 6.1 for computational evidence. Nevertheless, we focus on developing techniques for obtaining mixed-integer programming formulations of (HAP), which will allow us to leverage modern mixed-integer program solvers such as GUROBI and MOSEK. Then, in Section 4, we will use these techniques to find asymptotically optimal cyclic policies.
@@
$\max_{\boldsymbol{x},\boldsymbol{y},\boldsymbol{\rho}}\quad$ $\frac{1}{T}\sum_{i\in[N]}\sum_{t\in[T]}r_{i}y_{i}^{t}$
$\operatorname{s.t.}\quad$ $\boldsymbol{x}^{t}\in\{0,1\}^{N}\cap\mathcal{X}\text{ and }\boldsymbol{x}\in\mathcal{P}$ $\operatorname{for}\,t\in[T]$
@@
$\alpha_{i}(z^{1},\ldots,z^{M}):=\exp\biggl{(}\beta_{i}^{0}+\sum_{m\in[M]}\beta^{m}_{i}\cdot z^{m}\biggr{)}\qquad\operatorname{for}\;(z^{1},\ldots,z^{M})\in\{0,1\}^{M}.$
For a given assortment plan $\boldsymbol{x}:=(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{T})$, we interpret $\alpha_{i}(x_{i}^{t-1},\ldots,x_{i}^{t-M})$ as the attractiveness of product $i$ in period $t$. Therefore, we will henceforth refer to $\alpha_{i}(\cdot)$ as the attraction value function. To handle the discrete nature of the attraction value function, we will use the notion of a continuous extension. Given a function $f:\{0,1\}^{n}\to\mathbbm{R}$, a function $g:[0,1]^{n}\to\mathbbm{R}$ is called a continuous extension of $f(\cdot)$ if $g(\cdot)$ is continuous and $g(\boldsymbol{z})=f(\boldsymbol{z})$ for every $\boldsymbol{z}\in\{0,1\}^{n}$. In addition, if an extension function $g(\cdot)$ is convex (resp. concave), we say that it is a convex (resp. concave) extension. It is common for a discrete function to admit many convex (resp. concave) extensions. A natural idea is to use the tightest convex/concave extension. Since the discrete domain $\{0,1\}^{n}$ is the vertex set of the continuous domain $[0,1]^{n}$, the pointwise largest convex (resp. smallest concave) extension of $f(\cdot)$ is the convex (resp. concave) envelope, denoted as $\operatorname{conv}(f)(\cdot)$ (resp. $\operatorname{conc}(f)(\cdot)$), of $f(\cdot)$ over the box $[0,1]^{n}$ (see Tawarmalani and Sahinidis 2002, Theorem 6).
3.1.1 A Lifted Representation of (Choice)
In this subsection, we focus on modeling constraint (Choice). The idea is to lift the original nonconvex constraint (Choice) into a higher-dimensional space where the lifted constraint is dramatically simpler to represent than the original constraint. Specifically, this will allow us to invoke prevalent techniques, such as the McCormick envelope (McCormick 1976), the perspective formulation (Günlük and Linderoth 2010), and the Lovász extension (Lovász 1983), to model the lifted nonconvex structures.
@@ 3.1.1 A Lifted Representation of (Choice)
where the first constraint can be linearized using the McCormick envelope McCormick (1976) and the second constraint is convex, since the function $\lambda f(\boldsymbol{u}/\lambda)$ is the perspective function of the convex function $f(\cdot)$ (Rockafellar 1970). Formally, given a function $f:D\to\mathbbm{R}$, where $D$ is a bounded subset of $\mathbbm{R}^{n}$, the perspective function of $f$, denoted as $\operatorname{pers}(f)$, is defined as
@@
Constraint (2a) lifts the nonlinear scaling factor $\rho^{t}x_{i}^{t}$ to $\gamma_{i}^{t}$; constraint (2b) lifts the scaled historical assortment decision $\gamma_{i}^{t}x_{i}^{t-m}$ to $z_{im}^{t}$; and constraint (2c) perspectifies the attraction value function in the space of the newly introduced variables $\gamma_{i}^{t}$ and $\boldsymbol{z}_{i}^{t}$.
@@
$z_{im}^{t}$ $\leq\min\bigl{\{}\gamma_{i}^{t},\rho^{t}_{U}x^{t-m}_{i}\bigr{\}}$ $\quad\operatorname{for}\,m\in[M]$ (4)
$z_{im}^{t}$ $\geq\max\bigl{\{}0,\ \rho^{t}_{U}x^{t-m}_{i}+\gamma^{t}_{i}-\rho^{t}_{U}\bigr{\}}$ $\quad\operatorname{for}\,m\in[M].$
@@
This is a convex inequality since $\tilde{\alpha}_{i}(\cdot)$ is a convex function and so is its perspective function (Rockafellar 1970). Moreover, this constraint can be represented as an exponential cone constraint (Ben-Tal and Nemirovski 2001) (see more details in the ecompanion EC.3.2):
@@
Next, we show that an explicit linear description of this constraint can be obtained from perspectifying a variant of the Lovász extension of $\alpha_{i}(\cdot)$ (Lovász 1983).
@@
The Concave Envelope of $\alpha_{i}(\cdot)$ Our construction is based on the Lovász extension of a set function. The Lovász extension is an extension of a function $f$ defined on $\{0,1\}^{n}$ to a function $\hat{f}$ defined on $[0,1]^{n}$. For each subset $S$ of $[n]$, let $\chi^{S}\in\{0,1\}^{n}$ be its indicator vector—that is, the $i^{\text{th}}$ coordinate of $\chi^{S}$ is $1$ if and only if $i\in S$. Observe that every vector $\boldsymbol{x}\in[0,1]^{n}$ can be expressed uniquely as $\boldsymbol{x}=\lambda_{0}\chi^{T_{0}}+\lambda_{1}\chi^{T_{1}}+\cdots+\lambda_{n}\chi^{T_{n}}$, where $\lambda_{k}\geq 0$ for $k=0,1,\ldots,n$ and $\sum_{k=0}^{n}\lambda_{k}=1$, and $\emptyset=T_{0}\subseteq T_{1}\subseteq\cdots\subseteq T_{n}=[n]$. Thus,
$\hat{f}(\boldsymbol{x})=\lambda_{0}f(\chi^{T_{0}})+\lambda_{1}f(\chi^{T_{1}})+\cdots+\lambda_{n}f(\chi^{T_{n}})$
@@
is a well-defined extension of the function $f$ (called the Lovász extension of $f$) on the continuous domain $[0,1]^{n}$. The Lovász extension serves as a bridge between supermodularity and concavity: $f$ is supermodular if and only if its Lovász extension $\hat{f}$ is concave (Lovász 1983). Moreover, when the Lovász extension $\hat{f}$ is concave, it is expressible as the pointwise minimum of affine functions (Tawarmalani et al. 2013).
When the history-dependent effects are negative, the attraction value function $\alpha_{i}(\cdot)$ is supermodular, and its concave envelope is given by its Lovász extension (Tawarmalani et al. 2013). The key idea behind the closed-form description of the Lovász extension is to interpolate the attraction value function affinely over the nested historical assortments $\chi^{T_{0}},\chi^{T_{1}},\ldots,\chi^{T_{M}}$, where $T_{0}\subseteq T_{1}\subseteq\cdots\subseteq T_{M}$. These nested assortments can be thought of as a path through the set of all possible past assortments $\{0,1\}^{M}$ going from $(0,0,\ldots,0)$ to $(1,1,\ldots,1)$, with each step along one coordinate direction. For example, when the memory length is two, the nested assortments $(0,0),(1,0)$, and $(1,1)$ correspond to the blue path in Figure 1(a). This path starts at $(0,0)$, moves along the first coordinate, then the second, and finally reaches $(1,1)$. To treat the general case of mixed effects, we modify the nested assortments—or equivalently the path—as follows: Instead of moving from $(0,0,\ldots,0)$ to $(1,1,\ldots,1)$, we start at $\chi^{I}$ and end at $\chi^{M\setminus I}$, where $I$ denotes the indices of positive effects. For instance, consider mixed effects with $\beta_{i}^{1}>0$ and $\beta_{i}^{2}\leq 0$. The blue path in Figure 1(b), starting from $(1,0)$ and ending at $(0,1)$, corresponds to the past assortments that we will use in this case.
Figure 1: Illustration of switched nested assortments under two different history-dependent effects
(a) Satiation effects: $\beta_{i}^{1}\leq 0,\beta_{i}^{2}\leq 0$
$(0,0)$$(1,0)$$(1,1)$$(0,1)$$x_{i}^{t-1}$$x_{i}^{t-2}$
(b) Mixed effects: $\beta_{i}^{1}>0,\beta_{i}^{2}\leq 0$
$(0,0)$$(1,0)$$(0,1)$$(0,1)$$1-x_{i}^{t-1}$$x_{i}^{t-2}$
Now, we formally define the historical assortments used in our construction. For each product $i\in[N]$, let $I_{i}$ denote the indices of positive effects, that is, $I_{i}:=\{m\mid\beta_{i}^{m}>0\}$. Let $\Omega$ denote all permutations of past periods $[M]$. For each product $i\in[N]$ and a permutation $\sigma\in\Omega$, we define $M+1$ historical assortments, denoted as $\boldsymbol{h}^{\sigma}_{i,0},\boldsymbol{h}^{\sigma}_{i,1},\ldots,\boldsymbol{h}^{\sigma}_{i,M}$, as follows: $\boldsymbol{h}^{\sigma}_{i,0}=\chi^{I_{i}}$ and for each $k\in[M]$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{h}^{\sigma}_{i,k}=\begin{cases}\boldsymbol{h}^{\sigma}_{i,k-1}+\boldsymbol{e}_{\sigma(k)}&\text{ if }\sigma(k)\notin I_{i}\\ \boldsymbol{h}^{\sigma}_{i,k-1}-\boldsymbol{e}_{\sigma(k)}&\text{ if }\sigma(k)\in I_{i},\end{cases}}$ (SwitchNested)
With this definition, we are ready to describe the concave envelope of $\alpha_{i}(\cdot)$ which, after perspectification, yields an explicit linear description for the constraint (6).
Proposition 3.1
Constraint (6) is equivalent to the following system of linear inequalities.
@@
$y^{t}_{i}\leq\alpha_{i}(\boldsymbol{h}^{\sigma}_{i,0})(\gamma^{t}_{i}-\tilde{z}_{i\sigma(1)}^{t})+\sum_{k\in[M]}\alpha_{i}(\boldsymbol{h}^{\sigma}_{i,k})(\tilde{z}_{i\sigma(k)}^{t}-\tilde{z}_{i\sigma(k+1)}^{t})\qquad\text{ for }\sigma\in\Omega\$ (7)
where $\tilde{z}_{i\sigma(M+1)}^{t}=0$, $\tilde{z}^{t}_{i\sigma(k)}=z^{t}_{i\sigma(k)}$ if $\sigma(k)\notin I_{i}$, and $\tilde{z}^{t}_{i\sigma(k)}=\gamma_{i}^{t}-z^{t}_{i\sigma(k)}$ if $\sigma(k)\in I_{i}$.
3.1.2 The Final Formulation